# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 4.2.2. Data Poisoning Attack
Data poisoning attacks refer to intentionally manipulating the training data of an AI model to disrupt its decision-making processes. Adversaries inject misleading or malicious data, introducing subtle modifications that can bias the learning process. This manipulation leads to incorrect outputs and faulty decision-making by the AI model (poision1, ). Manipulating the behavior of these deep learning systems according to the attacker’s intentions can be achieved by poisoning the training data. Several studies demonstrated that adversaries can insert poison examples into datasets used to train language models (LMs) (wan2023poisoning, ), (kurita2020weight, ). Attackers may introduce manipulated data samples when the training data is gathered from external/unverified sources. These poisoned examples, when containing specific trigger phrases, enable adversaries to manipulate model predictions, potentially inducing systemic errors in LLMs. Trojan attacks can be achieved through data poisoning, where the malicious data is injected into the training to create a hidden vulnerability or a ’Trojan trigger’ in the trained model, causing abnormal model behaviors when activated by specific triggers. Zhang et al. introduced TROJANLM, a trojan attack variant where specially crafted LMs induce predictable malfunctions in host NLP systems (zhang2021trojaning, ). Traditional poisoning attacks involve directly injecting insecure code into the training data. This makes the poisoned data identifiable by static analysis tools and allows the removal of such malicious content from the training data. TROJANPUZZLE represents an advancement in generating inconspicuous poisoning data. It ensures that the model suggests the complete payload during the code generation outside docstrings by removing suspicious portions of the payload in the poisoned data (aghakhani2023trojanpuzzle, ). While several baseline defense techniques, such as training sample filtering, reorganization, and rephrasing, have been proposed in (jain2023baseline, )), handling adversarial inputs remains challenging for LLMs. Wang et al. conduct a thorough evaluation of the efficacy of ChatGPT from the adversarial and OOD aspects (wang2023robustness, ). Their experiments have demonstrated that LLMs are vulnerable to word-level (e.g., typo) and sentence-level (e.g., distraction) adversarial inputs. Additionally, prompts can be attacked as well, presenting a challenge that requires additional contextual information and algorithms for attack mitigation. This is currently a complex and challenging problem due to the high sensitivity of LLMs to prompts (maus2023adversarial, ).
@@ 5. Privacy Attacks of LLMs
Privacy risks in LLMs arise from their inherent capacity to process and generate text based on extensive and diverse training datasets. These models, like GPT-3, may inadvertently capture and reproduce sensitive information that exists in training data, potentially posing privacy concerns during the text generation process. Issues such as unintentional data memorization, data leakage, and the potential disclosure of confidential or PII are key challenges (pan2020privacy, ). Fine-tuning LLMs for specific tasks introduces additional privacy considerations. Making a balance between the utility of these powerful language models and the imperative to protect user privacy is very crucial for assuring the reliable and ethical use of LLMs in various applications. In Figure 6, we show the categories of privacy attacks in LLMs with some examples.
Figure 6. Privacy Attacks in LLMs and Examples.
@@ 5.1. Gradient Leakage Attack
Deep learning models are often trained using optimization algorithms that involve gradients. Gradients represent the direction of the steepest increase in a function that helps to optimize model parameters during training to minimize the loss function. If attackers can access or infer these gradients or the gradient information, they may obtain access to the model or even compromise its privacy and safety, e.g., reconstructing private training data (gradAttck, ). Sensitive information can be extracted by analyzing gradients during the training or manipulating the training data. Several studies (wei2020framework, ), (zhu2019deep, ), (geiping2020inverting, ), (das2023privacy, ) have shown private training samples can be successfully reconstructed from the deep learning model using gradients with high reconstruction accuracy under the federated learning environment. However, those algorithms work mostly for image datasets. Very few studies have studied the gradient leakage attacks for LMs. These are small language models trained over far fewer parameters than LLMs, e.g., TinyBERT (jiao2019tinybert, ). Deng et al. introduced a universal gradient attack on Transformer-based language models in the NLP domain. They call it TAG, which can reconstruct the private training samples, $(X,Y)$, from Transformer-based LMs (deng2021tag, ).
The TAG adversary acquires gradients $\nabla W$ from a participant in a distributed learning system, then updates randomly initialized dummy data $(X_{0},Y_{0})$ through a comparison of the difference between acquired gradients $\nabla W$ (from clients) and the adversary’s gradients $\nabla W_{0}$. This is achieved by leveraging a loss function, e.g., L1-norm (Manhattan distance) or L2-norm (Euclidean distance), along with a coefficient parameter $\alpha$. Eventually, the adversary can access the client’s private information by recovering the private training data $(X,Y)$.
Compared to the existing method (zhu2019deep, ), TAG operates on models that are initialized with more realistic and pretrained weight distributions. According to (deng2021tag, ), TAG can effectively reconstruct up to 88.9% tokens with 0.93 cosine similarity to the private training data.
The LAMP attack is a technique designed for recovering input text from gradients in a Federated Leaning (FL) environment proposed by Balunovic et al. (balunovic2022lamp, ).
It utilizes an auxiliary language model to assist in guiding the search process toward generating natural text (balunovic2022lamp, ). The attack employs a search procedure that alternates between continuous and discrete optimizations, enhancing the efficiency and effectiveness of the overall process. According to their experiments, LAMP performs better than previous methods (deng2021tag, ), reconstructing 5 times more bi-grams and achieving an average of 23% longer subsequences (balunovic2022lamp, ). Additionally, their proposed approach is the first to successfully restore input data from batch sizes greater than 1. These aforementioned attack models primarily targeted the language models (zhu2019deep, ), (deng2021tag, ), but they can also be potentially applied to LLMs. It is imperative to conduct additional comprehensive research and analysis to assess the impacts of these attacks on LLMs and identify effective mitigation strategies.
@@ 5.2. Membership Inference Attack
The primary goal of a Membership Inference Attack (MIA) is to determine if a data sample has been included in an ML model’s training data (erkin2009privacy, ), (yang2022holistic, ). The attackers can execute MIAs even in the absence of direct access to the underlying ML model parameters, relying solely on the observation of its output (MIA1, ). Typically, such attacks take advantage of models’ tendency to overfit their training data, resulting in lower loss values for training samples (sablayrolles2019white, ). The LOSS attack is a straightforward baseline, which considers samples as training members if their loss values are less than a specified threshold (yeom2018privacy, ). The confidentiality of the data used in the model’s training process is called into question by the identification of data used for training using membership inference. These types of attacks pose serious privacy concerns, particularly in scenarios where the targeted model has undergone training on sensitive information, e.g., medical or financial data (truex2019demystifying, ). Shokri et al. first introduced MIA against ML Model (MIA-ML) (shokri2017membership, ). The attack model is trained through their proposed shadow training technique: First, several “shadow” models are constructed to mirror the behavior of the target model, where the training dataset is known, and thus so as is the ground truth about the membership. Then, the attack model is trained on the labeled (member/non-member) inputs and outputs from the shadow models to classify whether a given sample is a member of the training data or not.
Existing MIA against language models are mainly focused on text generation and downstream text classification tasks (song2019auditing, ), (shejwalkar2021membership, ). Xin et al. took the first initiative to perform a systematic audit of the privacy risks associated with Pretrained Language Models (PLMs) by focusing on the perspective of MIA (xin2022membership, ). They have shown how an adversary seeks to determine if a data sample belongs to the training data of PLMs in the practical and prevalent situation where downstream service providers often construct models derived from four different PLMs architectures (BERT, ALBERT, RoBERTa, XLNet). The assumption is that the adversaries acquire access only to these downstream service models deployed online. Additionally, they considered another more realistic scenario where no additional information about the target PLMs is available to the adversary other than the output, i.e., black-box setting. Most existing attacks in the literature rely on the fact that models often assign their training samples with higher probabilities than non-training instances. However, this approach tends to result in high false-positive rates as it overlooks the inherent complexity of a sample. For training the attack models (shokri2017membership, ), attacks of this type are based on a highly optimistic and arguably unrealistic assumption in many cases that an adversary knows the distribution of training data of the target model (ye2022enhanced, ).
@@
In applications, where the domain-specific (e.g., in the medical domain) publicly available data is not prevalent, these reference-based attacks are supposed to be ineffective in those cases (tramer2022considerations, ). Mattern et al. proposed a neighborhood attack that relies on the concept of using neighboring samples, generated through data augmentations like word replacements, as references for inferring membership, which aims to develop a metadata-free mechanism (mattern2023membership, ). Mireshghallah et al. reported that previous attacks on Masked Language Models (MLMs) ((kenton2019bert, ), (liu2019roberta, )) may have yielded inconclusive results due to their exclusive reliance on the loss of the target model on individual samples for the evaluation of how effectively the model memorized those samples (mireshghallah2022quantifying, ). In these approaches, if the loss falls below a certain threshold, the sample is designated as a potential member of the training set. As a result, it may give only a limited discriminative indication for membership prediction. Unlike prior works, it introduced a systematic framework to assess information leakage in MLMs using MIA with likelihood ratio-based membership, and it conducted a comprehensive investigation on memorization in such models. The attacks on conventionally non-probabilistic models become possible when MLMs are treated as probabilistic models over sequences. The attack method was evaluated on a collection of masked clinical language models and compared its performance against a baseline approach that completely relies on the loss of the target model, as established in previous work ((yeom2018privacy, ), (song2020information, )). Some works investigated MIA methods in language models on specific domains, e.g., clinical language models. Jagannatha et al. investigated the risks (jagannatha2021membership, ) of training-data leakage to estimate the empirical privacy leaks for model architectures such as BERT (devlin2018bert, ) and GPT-2 (radford2019language, ).
@@
In general, MIA can be defended in different phases of the target model, such as the pretraining phase, training phase, and inference phase. Various technologies to defend the LMs against MIA have been proposed, such as regularization, transfer learning, and information perturbation (hu2023defenses, ). Some MIA defense strategies have been proposed, for instance, information perturbation, to safeguard natural language models (kandpal2022deduplicating, ), however, it is not sufficient from all the perspectives of MIA in language models. Moreover, the aforementioned attack models mostly focused on language models. However, those can be applied to the LLMs as well. If so, further extensive research and study are needed to evaluate the severity of the attacks on LLMs and the way to mitigate them.

### v2

@@ 4.2.2. Data Poisoning Attack
Data poisoning attacks refer to intentionally manipulating the training data of an AI model to disrupt its decision-making processes. Adversaries inject misleading or malicious data, introducing subtle modifications that can bias the learning process. This manipulation leads to incorrect outputs and faulty decision-making by the AI model (Fox, 2023). Manipulating the behavior of the DNNs systems according to the attacker’s intentions can be achieved by poisoning the training data. Several studies demonstrated that adversaries can insert poisoned instances into datasets used to train/fine-tune LMs (Wan et al., 2023), (Kurita et al., 2020). Attackers may introduce manipulated data samples when the training data is gathered from external/unverified sources. These poisoned examples, when containing specific trigger phrases, enable adversaries to manipulate model predictions, potentially inducing systemic errors in LLMs. Trojan attacks can be achieved through data poisoning, where the malicious data is injected into the training to create a hidden vulnerability or a ’Trojan trigger’ in the trained model. It causes abnormal model behaviors when activated by specific triggers. A simple automated data poisoning pipeline named AutoPoision (Shu et al., 2023) was introduced by Shu et al. on OracleLM, which automatically modifies the clean instruction by initiating an adversarial context, e.g., using a particular keyword in the response. Zhang et al. introduced TROJANLM, a trojan attack variant where specially crafted LMs induce predictable malfunctions in the host NLP systems (Zhang et al., 2021b). Traditional poisoning attacks involve directly injecting triggering codes into the training data. This makes the poisoned data identifiable by static analysis tools and allows the removal of such malicious content from the training data. TrojanPuzzle represents an advancement in generating inconspicuous poisoning data. It ensures that the model suggests the complete payload during the code generation outside docstrings by removing suspicious portions of the payload in the poisoned data (Aghakhani et al., 2023). One of the early poisoning attacks on code completion models was proposed by Schuster et al. (Schuster et al., 2021). They proposed two types of attacks: model poisoning and data poisoning. The attacker directly manipulates the autocompleter by fine-tuning it on carefully crafted files to perform model poisoning attacks. On the other hand, in data poisoning attacks, the attacker introduces these crafted files into open-source code repositories, which are used to train the autocompleter. The generation of accurate responses through LLMs is highly dependent on the quality and specificity of the instructions provided by its users, as these instructions guide the model’s responses. Consequently, an adversary could easily exploit this instruction-following behavior of LLMs to make it more susceptible to poisoning attacks. CodeBreaker proposed by Yan et al. utilizes poisoned data during the fine-tuning stage of the code completion model (Yan et al., 2024a) without impacting core functionalities. It can also deceive the vanilla vulnerability detectors of target CodeGen models (Mansfield et al., 2022), such as CodeGen-NL, CodeGen-Multi, and CodeGen-Mono. Recently, Retrieval Augmented Generation (RAG) has been used to enhance the domain-specific capabilities of LLMs by connecting the model to external data sources, such as an organization’s internal knowledge base, beyond its training data (Fan et al., 2024). It enhances the capability of LLMs by retrieving relevant data to perform domain-specific tasks more accurately (Zhao et al., 2024). Poisoning attacks (e.g., AgentPoison (Chen et al., 2024b)) can also be conducted on RAG-based LLM agents by injecting a small number of poisoned samples into the RAG knowledge base. Several studies (Liang et al., 2024) illustrated poisoning attacks on vision language models (VLM). For example, in Shadowcast (Xu et al., 2024), poisoning is performed (with a minimal number of poisoned samples, as less as 50) so that it becomes impossible to differentiate the poisoned model from a benign model visually with the same text. The attack works in two perspectives in the inference stage: label attacks (which cause incorrect prediction) and persuasion attacks (which cause misinformation or incorrect judgment). The empirical study demonstrated that the method is highly effective and adaptable across various VLM architectures, even in black-box environments. A prompt-specific poisoning attack named NightShade (Shan et al., 2023) can control the output of a prompt in text-to-image generative models with as few as 100 poisoned training instances. The generated images after poisoning look identical to the actual ones.
Basic data poisoning attacks (e.g., You Auto-complete Me (Schuster et al., 2021), TROJANLM (Zhang et al., 2021b), and AutoPoison (Shu et al., 2023)) can be mitigated by using existing defense techniques such as static analysis, fine-pruning (Liu et al., 2018a), data curation (Yan et al., 2024a), and STRIP (Gao et al., 2019). Advanced methods, such as TrojanPuzzule (Aghakhani et al., 2023), rely on carefully crafted triggers for success. Arbitrarily designed triggers may not be able to activate backdoors. Additionally, some attacks, such as AgentPoison (Chen et al., 2024b), are based on assumptions that the attacker has black-box access to the models. Such assumptions may not be realistic in real-world scenarios. Furthermore, several methods, e.g., NightShade (Shan et al., 2023) and AutoPoision (Shu et al., 2023), were evaluated on LMs or small-scale LLMs. Thus, their performance on large-scale LLMs with hundreds of billions of parameters still remains uncertain. Though several baseline defense techniques, such as training sample filtering, reorganization, and rephrasing, have been proposed in (Jain et al., 2023), handling adversarial inputs remains challenging for LLMs. Wang et al. conduct a thorough evaluation of the efficacy of ChatGPT from the adversarial and OOD aspects (Wang et al., 2023a). Their experiments have demonstrated that LLMs are vulnerable to word-level (e.g., typo) and sentence-level (e.g., distraction) adversarial inputs. Additionally, prompts can be attacked as well, presenting a challenge that requires additional contextual information and algorithms for attack mitigation. This is currently a complex and challenging problem due to the high sensitivity of LLMs to prompts (Maus et al., 2023). Apart from that, query rephrasing (Kumar et al., 2023), fine-pruning (Liu et al., 2018a), and data curation (Yan et al., 2024a) can also be utilized as potential defense techniques for data poisoning attacks.
Privay Attacks in LLMsAttacksGradint leakageMembershipInferenceAttacksPIILeakageAttacksExamples:1. TAG* (Deng et al., 2021)2. LAMP (Balunovic et al., 2022)3. GDBA (Guo et al., 2021)Examples:1. MIA-ML* (Shokri et al., 2017)2. MIA on PLM (Xin et al., 2022)3. PREMIA (Feng et al., 2024b)4. SaMIA (Kaneko et al., 2024)Examples:1. TAB* (Inan et al., 2021b)2. ProPILE(Kim et al., 2024)3. PII Compass (Nakka et al., 2024)4. Memorization (Carlini et al., 2021)5. KART* (Nakamura et al., 2020)*Some attacks were performed only on language models
Figure 8. Privacy Attacks in LLMs and Examples.
@@ 5. Privacy Attacks of LLMs
Privacy risks in LLMs arise from their inherent capacity to process and generate text based on extensive and diverse training datasets. These models, like GPT-3, may inadvertently capture and reproduce sensitive information that exists in training data, potentially posing privacy concerns during the text generation process. Issues such as unintentional data memorization, data leakage, and the potential disclosure of confidential information or PII are key challenges (Pan et al., 2020). Fine-tuning LLMs for specific tasks introduces additional privacy considerations. Making a balance between the utility of these powerful LMs and the imperative to protect user privacy is very crucial for assuring the reliable and ethical use of LLMs in various applications. In Figure 8, we show the categories of privacy attacks in LLMs with some examples. Additionally, in Table 4, we briefly mention the LLM privacy attack categories, source code, their characteristics, limitations and potential mitigation techniques, in corresponding subsections, we further discuss it in detail.
@@ 5.1. Gradient Leakage Attack
Deep learning models are often trained using optimization algorithms that involve gradients. Gradients represent the direction of the steepest increase in a function that helps to optimize model parameters during training to minimize the loss function. If attackers can access or infer these gradients or the gradient information, they may obtain access to the model or even compromise its privacy and safety, e.g., reconstructing private training data (Haldar, 2023). Sensitive information can be extracted by analyzing gradients during the training or manipulating the training data. Several studies (Wei et al., 2020), (Zhu et al., 2019), (Geiping et al., 2020), (Das et al., 2023) have shown private training samples can be successfully reconstructed from the deep learning model using gradients with high reconstruction accuracy under the Federated Learning (FL) environment. However, those algorithms work mostly for image datasets. Very few studies have investigated the gradient leakage attacks for LMs. These are small LMs trained over far fewer parameters than LLMs, e.g., TinyBERT (Jiao et al., 2019).
One of the early explorations of gradient leakage attacks against LMs is gradient-based distributional attack (GBDA) which performs gradient-based text attacks on Transformers (Guo and Kiela., 2021). It leverages parameterized adversarial distribution to enable gradient optimization to perform efficient gradient leakage attacks instead of using a single adversarial instance. This attack is effective for different LMs on distinct tasks. Deng et al. introduced a universal gradient attack on Transformer-based LMs in the NLP domain, named TAG. TAG can reconstruct the private training samples, $(X,Y)$, from Transformer-based LMs (Deng et al., 2021). The TAG adversary acquires gradients $\nabla W$ from a participant in a distributed learning system, then updates randomly initialized dummy data $(X_{0},Y_{0})$ through a comparison of the difference between acquired gradients $\nabla W$ (from clients) and the adversary’s gradients $\nabla W_{0}$. This is achieved by leveraging a loss function, e.g., L1-norm (Manhattan distance) or L2-norm (Euclidean distance), along with a coefficient parameter $\alpha$. Eventually, the adversary can access the client’s private information by recovering the private training data $(X,Y)$. Compared to the existing method (Zhu et al., 2019), TAG operates on models that are initialized with more realistic and pre-trained weight distributions. According to Deng et al., TAG can effectively reconstruct up to 88.9% tokens with 0.93 cosine similarity in token embeddings to the private training data (Deng et al., 2021). The LAMP attack is a technique designed for recovering input text from gradients in an Federated Leaning (FL) environment proposed by Balunovic et al. (Balunovic et al., 2022). It utilizes an auxiliary LM to guide the search process toward generating natural text (Balunovic et al., 2022). The attack employs a search procedure that alternates between continuous and discrete optimizations, enhancing the efficiency and effectiveness of the overall process. According to their experiments, LAMP performs better than previous methods (Deng et al., 2021), reconstructing 5$\times$ more bi-grams and achieving an average of 23% longer subsequences (Balunovic et al., 2022). Additionally, their proposed approach was the first to successfully restore input data from batch sizes greater than 1. These aforementioned attack models primarily targeted the LMs (Zhu et al., 2019), (Deng et al., 2021), however they can also be potentially applied to LLMs. Another essential condition of the gradient-based attack methods is they often need white-box access (e.g., gradients) to the model for performing attacks. Even with gradient access, attackers are often limited to reconstructing the training data, sometimes only at the token level (Guo et al., 2021), rather than the training labels (Balunovic et al., 2022).
To defend against gradient leakage attacks, noise perturbation to the gradients (Zhu et al., 2019) and differential privacy (DP) (Abadi et al., 2016) are very popular methods in the computer vision domain. Moreover, successfully executing an attack against a well-trained model often requires specific settings, e.g., loss function and optimizer (Geiping et al., 2020). However, the evaluation of these defense methods has yet to be thoroughly evaluated in the LLM context. Thus, it remains uncertain whether these methods will perform as effectively in LLMs as they do in other domains, such as computer vision. In order to overcome these shortcomings, it is imperative to conduct additional comprehensive research and analysis to assess the impacts of these attacks on LLMs and identify effective mitigation strategies.
Category Name
Source
Code
Characteristics Limitation
Potential
Mitigation
Gradient Leakage Attacks TAG (Deng et al., 2021) GitHub
- Performs gradient attack to recover the local
training data on Transformer-based language model.
- Impact: Moderate
- Adversary requires access to the
gradient of the Transformer model.
- Noise Perturbation (Zhu et al., 2019).
- Differential Privacy (Abadi et al., 2016).
LAMP (Balunovic et al., 2022) GitHub
- Recovers training data from gradients in FL
environment using guided search with auxiliary LM.
- Impact: High
- Does not deal with reconstructing
training labels.
- Differential Privacy (Abadi et al., 2016)
GDBA (Guo et al., 2021) GitHub
- Introduced parameterized adversarial distribution
to perform gradient-based text-attack.
- The attack is transferable to other LMs and
application for varieties of tasks.
- Impact: Moderate
- The method sis only able to recover
tokens.
- Robust optimization
in training so that
adversarial perturbation
has minimum impact (Zhang et al., 2021a).
Membership Inference Attacks
MIA
on
PLMs
(Xin et al., 2022)
N/A
- Attacker collects a small number of training data
of the model which the PLMs have pre-trained
on and trains the attack model with these instances.
- Designs a binary classifier (member or non-
member), leads to MIA.
- Impact: Moderate
- Attack model stands on an unrealistic
assumption of having access to some
sample of training data.
- Adversarial regularization
(Nasr et al., 2018).
- MemGuard (Jia et al., 2019).
PREMIA (Feng et al., 2024b) N/A
- Performs MIA based on preferred and non-
preferred responses against specific prompts in LLMs.
- Based on the response of the model the attacker
determines the membership of the training set.
- Impact: Moderate
- Attack was fully evaluated on open-
source LLMs. The performance might
not be same on close-source model
such as GPT-3.
- DP-SGD (Abadi et al., 2016)
- Model pruning(Han et al., 2015).
- Knowledge distillation (Hinton, 2015)
SaMIA (Kaneko et al., 2024) GitHub
- Introduces sampling-based pseudo likelihood (SPL)
method for MIA based on the n-gram similarity
between pre-fix (part of generated text) and
reference text (remaining part of the generated text).
- Impact: Moderate
- Attacks performed on LLMs which
training data is known.
- Performance might degrade on the
unknown data on which model has
been trained.
- To be studied.
PII Leakage Attacks TAB (Inan et al., 2021b) N/A
- An evaluation technique if the PII can be leaked
under any leaked under any threat model.
- Impact: Low
- The proposed method was evaluated
on RNN and Transformer-based language
model, did not demonstrate applicability
on LLM or other model architectures.
- DP-SGD (Abadi et al., 2016)
- API Hardening (Inan et al., 2021b)
PII
Compass (Nakka et al., 2024)
N/A
- Proposed PII leakage method by prepending
hand-crafted template with true prefix subject
to different data in the adversary set.
- Impact: Low
- Limited to extracting only phone number
due to the lack of publicly available PII
entries such as SSN.
- Adversarial Training (Nakka et al., 2024)
- PII masking (Mansfield et al., 2022)
- DP-SGD (Abadi et al., 2016)
ProPILE (Kim et al., 2024) N/A
- Introduced as a tool for risk assessment of PII
leakage from open pre-trained transformer
language model.
- Based on the linkability of the given information
of asked query evaluates the risk of PII exposure.
- Impact: Moderate
- The performance highly depends on the
evaluation that is available on open-
source datasets.
- The evaluation dataset might contain
noisy or misleading information.
- Cleaning PII with structured
patterns with regular
expression (Kim et al., 2024)
Memorization (Carlini et al., 2021) GitHub
- Demonstrated the PII leakage attack due to
memorization in LLMs, i.e., GPT-2.
- The proposed attack efficiently extracts training
samples using certain prefixes by black-box query
access
- Impact: Low
- The number of successful extraction
of PII are extremely low.
- The attack was evaluated on GPT-2,
it may not perform well on the latest LLms,
e.g., Llama, GPT-4.
- Training Data curation(Continella et al., 2017).
- DP-SGD (Abadi et al., 2016)
- Prompt-tuning (Ozdayi et al., 2023)
KART (Nakamura et al., 2020) GitHub
- Defines four factors that contributes to a privacy
leakage in PLMs, i.e., attacker’s prior knowledge
about the model, the target information to be leaked,
resources to attack, and the availability of the
target information in pre-trained data. (Kim et al., 2024)
- Impact: Low
- It did not provide any universal attack
methods in which all the factors contribut
ed to execute a successful attack.
- The impact on attack performance after
applying DP was not evaluated.
- Differential Privacy (Abadi et al., 2016)
- Limiting attacker access to
training data.
Table 4. The categories of LLM privacy attacks, source code, their basic characteristics, limitations, and potential mitigation techniques. (This table is discussed in detail in the corresponding parts of Section 5 and 6)
@@ 5.2. Membership Inference Attack
The primary goal of a Membership Inference Attack (MIA) is to determine if a data sample has been included in an ML model’s training data (Erkin et al., 2009), (Yang, 2022). The attackers can execute MIAs even in the absence of direct access to the underlying ML model parameters, relying solely on the observation of its output (Dickson, 2021). Typically, such attacks take advantage of models’ tendency to overfit their training data, resulting in lower loss values for training samples (Sablayrolles et al., 2019). The LOSS attack is a straightforward baseline where the basic idea is if their loss values are less than a specified threshold, then it considers samples as training members (Yeom et al., 2018). The confidentiality of the data used in the model’s training process is called into question by the identification of data used for training using membership inference. These type of attacks poses serious privacy concerns, particularly in scenarios where the targeted model has undergone training on sensitive information, e.g., medical or financial data (Truex et al., 2019). Shokri et al. first introduced MIA against ML Model (MIA-ML) (Shokri et al., 2017). The attack model is trained through their proposed shadow training technique: First, several “shadow” models are constructed to mirror the behavior of the target model, where the training dataset is known, and thus so is the ground truth about the membership. Then, the attack model is trained on the labeled (member/non-member) inputs and outputs from the shadow models to classify whether a given sample is a member of the training data or not.
Following the formal description of Shokri et al. (Shokri et al., 2017), given an attack target model $f_{\text{target}}()$, its private training dataset $D^{\text{train}}_{\text{target}}$ contains labeled training samples $(x^{i},y^{i})_{\text{target}}\in D^{\text{train}}_{\text{target}}$. Here, $x^{i}_{\text{target}}$ represents the model input, and $y^{i}_{\text{target}}$ is the ground truth label, taking a value from a set of $c_{\text{target}}$ classes. The target model will output a probability vector $Y^{i}$ of size $c_{\text{target}}$ as the prediction on input $x^{i}$. $f_{\text{attack}}()$ is the attack model, which is a binary classifier to determine whether a given sample $(x,y)$ is in the private training dataset (“in”) or not (“out”). It is challenging to distinguish between members and non-members in the private training dataset. Moreover, this task becomes increasingly difficult when the attacker has limited information about the internal parameters of the target model and can only access it through public APIs with a limited number of queries.
MIA-ML (Shokri et al., 2017) leverages shadow models to implement the attack model. Each shadow model $f^{i}_{shadow}()$ will be trained on a dataset $D^{\text{train}}_{{\text{shadow}_{i}}}$, which is similar in terms of format and distribution to, but disjoint from the private training dataset $D^{\text{train}}_{\text{target}}$, i.e., $D^{\text{train}}_{{\text{shadow}_{i}}}\cap D^{\text{train}}_{{\text{target}}}=\varnothing$. Once all $k$ shadow models are trained, the attack training set $D^{\text{train}}_{\text{attack}}$ can be generated by following (1) $\forall(x,y)\in D^{\text{train}}_{\text{shadow}_{i}}$, obtain the prediction vector (output) $Y=f^{i}_{\text{shadow}}(x)$ and include the record $(y,Y,\textit{in})$ in $D^{\text{train}}_{\text{attack}}$, and (2) query the shadow model with a test dataset $D^{\text{test}}_{{\text{shadow}_{i}}}$ disjoint from $D^{\text{train}}_{\text{shadow}_{i}}$, then $\forall(x,y)\in D^{\text{test}}_{\text{shadow}_{i}}$, generate the prediction vector (output) $Y=f^{i}_{\text{shadow}}(x)$ and add the record $(y,Y,\textit{out})$ to $D^{\text{train}}_{\text{attack}}$. $D^{\text{train}}_{\text{attack}}$ will be divided into $c_{\text{target}}$ partitions, where each partition is linked to a distinct class label. For each class label $y$, an individual attack model will be trained to predict the membership status “in” or “out” for a given input $(x,y)$.
@@
Existing privacy attacks against LMs for MIA are mainly focused on text generation and downstream text classification tasks (Song and Shmatikov, 2019), (Shejwalkar et al., 2021). Xin et al. took the first initiative to perform a systematic audit of the privacy risks associated with pre-trained language models (PLMs) by focusing on the perspective of MIA (Xin et al., 2022). They have shown how an adversary seeks to determine if a data sample belongs to the training data of PLMs in the practical and prevalent situation where downstream service providers often construct models derived from four different PLM architectures (BERT, ALBERT, RoBERTa, and XLNet). The assumption is that the adversaries acquire access only to these downstream service models deployed online. Additionally, they considered another more realistic scenario where no additional information about the target PLMs is available to the adversary other than the output, i.e., black-box setting. Most existing attacks in the literature rely on the fact that models often assign their training samples with higher probabilities than non-training instances. However, this approach tends to result in high false-positive rates as it overlooks the inherent complexity of a sample. For training the attack models (Shokri et al., 2017), attacks of this type are based on a highly optimistic and arguably unrealistic assumption in many cases that an adversary knows the distribution of training data of the target model (Ye et al., 2022). Mattern et al. proposed a neighborhood attack that relies on the concept of using neighboring samples, generated through data augmentations like word replacements, as references for inferring membership, which aims to develop a metadata-free mechanism (Mattern et al., 2023). Mireshghallah et al. reported that previous attacks on MLMs ((Kenton and Toutanova, 2019), (Liu et al., 2019)) may have yielded inconclusive results due to their exclusive reliance on the loss of the target model on individual samples for the evaluation of how effectively the model memorized those samples (Mireshghallah et al., 2022). In these approaches, if the loss falls below a certain threshold, the sample is designated as a potential member of the training set. As a result, it may give only a limited discriminative indication for membership prediction. Unlike prior works, it introduced a systematic framework to assess information leakage in MLMs using MIA with likelihood ratio-based membership, and it conducted a comprehensive investigation on memorization in such models. Conventionally, the attacks on non-probabilistic models become possible when MLMs are treated as probabilistic models over sequences. The attack method was evaluated on a collection of masked clinical language models and compared its performance against a baseline approach that completely relies on the loss of the target model, as established in previous works ((Yeom et al., 2018), (Song and Raghunathan, 2020)). Some works investigated MIA methods in LMs on specific domains, e.g., clinical language models. Jagannatha et al. investigated the risks (Jagannatha et al., 2021) of training-data leakage to estimate the empirical privacy leaks for model architectures such as BERT (Devlin et al., 2018) and GPT-2 (Radford et al., 2019). Kaneko et al. proposed a sampling-based pseudo-likelihood (SPL) method for MIA. They call it SaMIA (Kaneko et al., 2024). It can detect whether a given text is included in the training data of an LLM without requiring access to the model’s likelihood or loss values. SaMIA generates multiple text samples from the LLM during testing. It then calculates the n-gram overlap between the samples and the target text, using this overlap as a proxy to estimate the likelihood of the target text. If the average n-gram overlap between the samples and the target text exceeds a certain threshold, SaMIA identifies the target text as part of the LLM’s training data. Feng et al. introduced a novel reference-based attack framework, PREMIA (Feng et al., 2024b) (Preference data MIA) to analyze the vulnerability of using preference data in LLM alignment to membership inference attacks (MIAs). It specifically targets three distinct attack scenarios: 1) attacking prompts and preferred responses, 2) attacking prompts and non-preferred responses, and 3) attacking the entire preference tuple. Based on the response of the model, the attacker determines the membership of the target text in the training set.
Several limitations exist in MIA against LLMs. The concept of MIA fundamentally relies on the assumption that the attacker has white-box access to the model and training data. In practical scenarios, this assumption may be somewhat unrealistic. Additionally, SaMIA (Kaneko et al., 2024) was evaluated only on a single task, classification, and its performance on other tasks, such as translation or text generation, may not achieve similar effectiveness. Moreover, the method was tested on public training data WikiMIA (Shi et al., 2023a). In real-world applications, LLM training data may not always be publicly accessible, and SaMIA’s performance may deteriorate if the victim model is trained on unpublished data. On the other hand, PREMIA (Feng et al., 2024b) was evaluated on open-source LLMs, and its performance may not be consistent on closed-source models such as GPT-3.
@@
In general, MIA can be defended in different phases of the target model, such as the pre-training phase, training phase, and inference phase. Various technologies to defend the LMs against MIA have been proposed, such as regularization, transfer learning, and information perturbation (Hu et al., 2023b), (Kandpal et al., 2022). However, it is not sufficient from all the perspectives of MIA in LMs. The aforementioned attack models mostly focused on LMs. Potentially, those can be applied to the LLMs as well. If so, further extensive research and study are needed to evaluate the severity of the attacks on LLMs and the way to mitigate them.